# Certified algorithms for equilibrium states of local quantum Hamiltonians

### v2

@@ 1 Introduction
We further require that as $\ell\to\infty$, the interval $[\mathsf{p}^{\min}_{\ell},\mathsf{p}^{\max}_{\ell}]$ converges to the desired value $\mathsf{p}^{*}(H)$. Here, $\ell$ is a parameter that governs the runtime of the algorithm such that fast convergence in $\ell$ leads to an efficient algorithm. The above cited complexity results show e.g., that if $\mathsf{p}^{*}(H)$ is the ground energy of $H$, then for small $\ell$, the interval $[\mathsf{p}^{\min}_{\ell},\mathsf{p}^{\max}_{\ell}]$ has to be large for some $H$. However, for any input $H$ of interest, we can always run the algorithm and the returned information will be correct and certifiably so. As such, no a priori analysis of the convergence speed is needed in order to obtain rigorous approximations of the quantity of interest.
@@ Results
$[\langle O\rangle^{\min}_{\beta},\langle O\rangle^{\max}_{\beta}]:=\left\{\operatorname{tr}(\rho O):\rho\text{ is an equilibrium state at temperature $T=1/\beta$}\right\}.$ (2)
@@
We note that for finite systems, convergence happens after a finite number of steps, namely for $\ell$ such that $\Lambda_{\ell}=\Gamma$. While the convergence result becomes trivial in the finite case, the formulation of a convex program containing constraints for the Gibbs state still adds a novel tool for the numerical treatment of finite-sized quantum many-body systems. The more interesting case for our results is however that of infinite systems. Given a fixed computational budget, one can choose $\ell$ and run the corresponding program (3) and obtain a superset $[\mathsf{p}^{\min}_{\ell},\mathsf{p}^{\max}_{\ell}]$ of the target interval $[\langle O\rangle^{\min}_{\beta},\langle O\rangle^{\max}_{\beta}]$. In the cases where $\langle O\rangle^{\min}_{\beta}=\langle O\rangle^{\max}_{\beta}$, e.g., for finite systems at $T>0$, one can also run the program (3) with increasing values of $\ell$ until $\mathsf{p}^{\max}_{\ell}\leq\mathsf{p}^{\min}_{\ell}+\varepsilon$, where $\varepsilon$ is some desired accuracy $\varepsilon$. We note that, as previously mentioned, the problem of computing expectation values for equilibrium states is at least QMA-hard so we cannot hope to have fast convergence for all choices of $H$. But we stress that no a priori analysis of convergence speed is required to obtain some desired accuracy: as soon as we get $\mathsf{p}^{\max}_{\ell}\leq\mathsf{p}^{\min}_{\ell}+\varepsilon$, an additive $\varepsilon$ approximation is guaranteed for this instance. Numerical results illustrating this algorithm can be found in Section 4.
@@
$[\langle O\rangle^{\min,\mathrm{TI}}_{\beta},\langle O\rangle^{\max,\mathrm{TI}}_{\beta}]:=\left\{\operatorname{tr}(\rho O):\rho\text{ is a translation-invariant equilibrium state at temperature $T=1/\beta$}\right\}.$ (10)
@@
Note that this is a promise problem where the objective is to decide if all the equilibrium states have an expectation value per site $>a$ or all the equilibrium states have an expectation value per site $<a$.
@@
Consider a translation-invariant local Hamiltonian $H$ on $\Gamma=\mathbb{Z}^{D}$ with $D\leq 2$. Assume furthermore that $H$ is commuting, i.e., $[h_{X},h_{Y}]=0$ for all $X,Y\subset\mathbb{Z}^{D}$. For $D=1$ and $\beta_{1}=\infty$ or for $D=2$ and some $\beta_{1}>0$, we have for all $0\leq\beta<\beta_{1}$, and for any local observable $O$, $\langle O\rangle^{\min,\mathrm{TI}}_{\beta}=\langle O\rangle^{\max,\mathrm{TI}}_{\beta}=\langle O\rangle^{\mathrm{TI}}_{\beta}$ and, using the same notation as in Theorem 1.1
@@ 2 Preliminaries
where $h_{X}\in\mathcal{A}_{X}$ are the local interaction terms. We assume that the Hamiltonian $H$ is local, i.e., there exists a constant $r$ such that $h_{X}$ is nonzero only for $X$ having size at most $r$, for every site $x\in\Gamma$, |$|\{X\subset\Gamma:x\in X,h_{X}\neq 0\}|\leq r$, and that $\|h_{X}\|\leq 1$ for all $X$. In the case of the infinite lattice $\Gamma=\mathbb{Z}^{D}$, we say that the Hamiltonian is translation-invariant if $h_{X+x}=\tau_{x}(h_{X})$ for all $X\subset\mathbb{Z}^{D}$ and $x\in\mathbb{Z}^{D}$, where $\tau_{x}:\mathcal{A}_{X}\to\mathcal{A}_{X+x}$ is the translation operator by $x$. If $\Lambda$ is a finite subset of $\Gamma$, we let
@@
where the blocks correspond to the ground-space and its orthogonal complement, we can see that $C$ has nonpositive diagonal entries, but is also positive semidefinite and thereby 00. By the Schur complement lemma we can then conclude that $B=0$ and thereby $\rho$ is supported on the ground space.
@@
$h_{i,i+1}(s,s^{\prime})=\left\{\begin{array}[]{cc}0&\text{ if }s=s^{\prime}\in\{0,1\}\\ 2&\text{ if }(s,s^{\prime})\in\{(0,1),(1,0),(2,2),(2,1),(0,2),(1,2)\}\\ -1&\text{ if }s=2,s^{\prime}=0.\\ \end{array}\right.$ (30)
For the finite Hamiltonian $\tilde{H}_{\Lambda}$ with $\Lambda=\{-\ell,\dots,\ell\}$, the unique ground state is the sequence $(2,0,\ldots,0)$ with energy $-1$. This makes it also the ground state of the infinite system corresponding to the limit with open boundary condition. Importantly, note that for the observable $O=|0\rangle\!\langle 0|$ at site 00, we have that for any $\ell\geq 1$, its value in the ground state of $H_{\Lambda}$ is $1$. However, another valid ground state of the infinite chain is the all-1 state. For this ground state in the thermodynamic limit, the observable takes a value of 00. It is realized by adding a boundary term $h_{\ell+1}(s)=-2\delta_{1,s}$ as it lowers the energy of the all-1 configuration below the open boundary ground state.
@@ 3.1 Energy-entropy balance inequalities
Let $H$ be a local Hamiltonian. Then $\omega:\mathcal{A}\to\mathbb{C}$ is a $\beta$-KMS state if, and only if, it satisfies [^2] The function $x\log(x/y)$ is defined as equal to 0 if $x=0$ and +$\infty$+ ∞ if $x>0$ and $y=0$.
@@
Similarly to the example for ground-states in Remark 2.1, we demonstrate how different thermal states can arise from different choices of boundary conditions. Consider the 2D Ising model at some temperature $0<T<T_{c}$ (where $T_{c}$ is the critical temperature), and let for any $\ell\in\mathbb{N}$, $\rho_{\ell,\beta}$ be the finite Gibbs state on a region {,,}$\{-\ell,\dots,\ell\}^{2}$ with open boundary condition. Then by the spin-flip symmetry of the Ising Hamiltonian, we know that $\operatorname{tr}(Z_{x}\rho_{\ell,\beta})=0$ for any site $x$ in the region {,,}$\{-\ell,\dots,\ell\}^{2}$. Thus this means that for any limit $\omega_{\beta}$ of the finite Gibbs state with open boundary condition, we will have $\omega_{\beta}(Z_{x})=0$. It is known however that for the 2D Ising model below the critical temperature, there are two (extremal) infinite-volume Gibbs states $\omega_{\beta}^{+}$ and $\omega_{\beta}^{-}$ such that $\omega_{\beta}^{-}(Z_{x})<0<\omega_{\beta}^{+}(Z_{x})$, which can be reached by taking limits of finite Hamiltonians with fixed boundary condition, i.e., aligned spins on the boundary [48].
@@ 3.2 Convex relaxations
The EEB inequalities above can be used to obtain rigorous lower and upper bounds on the value of any observable on thermal states via convex optimization. Let $(\Lambda_{\ell})_{\ell\in\mathbb{N}}$ be an increasing sequence of subsets of $\Gamma$ such that the support of $O$ is contained in $\Lambda_{0}$, and such that $\Lambda_{\ell}\uparrow\Gamma$ as $\ell\to\infty$. To simplify the presentation, we will further assume that $\overline{\Lambda_{\ell-1}}\subset\Lambda_{\ell}$ for all $\ell$, where, as before, $\overline{\Lambda}=\Lambda\cup\partial^{ex}\Lambda$ (see (21)). For example, if $\Gamma=\mathbb{Z}^{D}$ and $H$ is a nearest-neighbor Hamiltonian and $O$ is supported on site 00, one can take $\Lambda_{\ell}=\{-\ell,\ldots,\ell\}^{D}$. For any $\ell$, let $\mathcal{A}_{\ell}\subset\mathcal{A}$ be the subalgebra of observables supported on $\Lambda_{\ell}$. Consider the optimization problem
@@ 3.3 Convergence
$[\langle O\rangle^{\min}_{\beta},\langle O\rangle^{\max}_{\beta}]=\{\omega(O):\omega\in\mathcal{G}_{\beta}\}.$ (48)
@@
so that we can write $[\tilde{\omega}(b_{s}^{*}b_{t})]_{s,t}=K^{*}[\tilde{\omega}(a_{i}^{*}a_{j})]_{i,j}K$ and similarly for $[\tilde{\omega}(b_{t}b_{s}^{*})]_{s,t}$. Hence we get
@@
For each $\ell$, let $\tilde{\omega}_{\ell}:\mathcal{A}_{\ell}\to\mathbb{C}$ be a feasible solution to (45) such that $\mathsf{p}^{\min}_{\ell}\leq\tilde{\omega}_{\ell}(O)\leq\mathsf{p}^{\min}_{\ell}+1/\ell$ (if the minimum value of the optimization problem is attained, we simply take $\tilde{\omega}_{\ell}$ the optimal solution which satisfies $\tilde{\omega}_{\ell}(O)=\mathsf{p}^{\min}_{\ell}$). We can extend $\tilde{\omega}_{\ell}$ to a state on the full $C^{*}$-algebra $\mathcal{A}$ [47, Proposition 2.3.24]. We now have a sequence of states $\tilde{\omega}_{\ell}$ on $\mathcal{A}$. Since the set of states on $\mathcal{A}$ is weak-* compact [47, Theorem 2.3.15] we can extract from $\tilde{\omega}_{\ell}$ a subsequence $\tilde{\omega}_{n(\ell)}$ that weak-* converges to a state $\tilde{\omega}$ on $\mathcal{A}$.
We now show that $\tilde{\omega}$ is a $\beta$-KMS state by showing that the conditions (2) hold. Let $a\in\mathcal{A}_{\text{loc}}$. Since $a$ is a local observable we have $a\in\mathcal{A}_{n(\ell)-1}$ for all large enough $\ell$. Hence this implies that
$\tilde{\omega}_{n(\ell)}([H,a])=0$ (51)
$\tilde{\omega}_{n(\ell)}(a^{*}a)\log\left(\frac{\tilde{\omega}_{n(\ell)}(a^{*}a)}{\tilde{\omega}_{n(\ell)}(aa^{*})}\right)\leq\beta\tilde{\omega}_{n(\ell)}(a^{*}[H,a])$
for all large enough $\ell$. By taking the limit $\ell\to\infty$ this shows that $\tilde{\omega}$ satisfies the conditions (2) and is thus a $\beta$-KMS state. Finally note that
$\tilde{\omega}(O)=\lim_{\ell}\tilde{\omega}_{n(\ell)}(O)\leq\lim_{\ell}\mathsf{p}^{\min}_{n(\ell)}+1/n(\ell)=\mathsf{p}^{\min}_{\infty}.$ (52)
Since $\mathsf{p}^{\min}_{\infty}\leq\langle O\rangle^{\min}_{\beta}\leq\tilde{\omega}(O)$ we necessarily get equality $\mathsf{p}^{\min}_{\infty}=\langle O\rangle^{\min}_{\beta}$. The same argument can be used to show that $\mathsf{p}^{\max}_{\infty}=\langle O\rangle^{\max}_{\beta}$. ∎
@@ 3.4 Decidability
$\min_{\|x\|_{2}^{2}=1}\langle x|[\tilde{\omega}(a_{i}^{*}a_{j})]_{ij}|x\rangle=\min_{\|x\|_{2}^{2}=1}\sum_{ij}\overline{x_{i}}x_{j}\omega(a_{i}^{*}a_{j})=\operatorname{tr}[\rho X^{*}X]\geq C_{\Lambda_{\ell}}\operatorname{tr}[X^{*}X]=C_{\Lambda_{\ell}}$ (61)
and similarly for $[\tilde{\omega}(a_{j}a_{i}^{*})]_{ij}$. We use this to strengthen our convex relaxation (45) by adding the contraints
@@
After this modification, since the moment matrix is bounded away from zero and upper bounded by 1, the eigenvalues in the argument of the logarithm for the matrix relative entropy are contained in an interval $[a,1/a]$ for some $a>0$. We now introduce a semidefinite approximation of $D_{op}$ from [17]. For any integer $m\geq 1$ let $h_{m}(x)=2^{m}\left(x^{1/2^{m}}-1\right)$ which satisfy $\log(x)\leq h_{m+1}(x)\leq h_{m}(x)$ and note that $h_{m}\to\log$ uniformly on any compact interval in $(0,\infty)$ since $0\leq h_{m}(x)-\log(x)\leq 2^{-m}[(x-1)^{2}+(x^{-1}-1)^{2}]$ for all $x>0$. Consider the noncommutative perspective [51, 52] of $-h_{m}$ which we denote by $D_{op}^{[m]}$:
@@
Then $D_{op}^{[m]}$ is jointly convex in $(X,Y)$, has a finite semidefinite programming formulation [17], and satisfies $D_{op}^{[m]}(X\|Y)\preceq D_{op}(X\|Y)$ since $-h_{m}\leq-\log$. By changing the function $D_{op}$ in the entropy constraint (45e) by $D_{op}^{[m]}$ we get a semidefinite program $(\mathsf{P}^{{}^{\prime}\min}_{\ell,m})$ with rational coefficients whose optimal value satisfies $\mathsf{p}^{{}^{\prime}\min}_{\ell,m}\leq\mathsf{p}^{{}^{\prime}\min}_{\ell}$ and analogously for the upper bounds.

### v3

@@ 1 Introduction
We further require that as $\ell\to\infty$, the interval $\mathsf{p}^{\min}_{\ell},\mathsf{p}^{\max}_{\ell}]$ converges to the desired value $\mathsf{p}^{*}(H)$. Here, $\ell$ is a parameter that governs the runtime of the algorithm such that fast convergence in $\ell$ leads to an efficient algorithm. The above cited complexity results show e.g., that if $\mathsf{p}^{*}(H)$ is the ground energy of $H$, then for small $\ell$, the interval $\mathsf{p}^{\min}_{\ell},\mathsf{p}^{\max}_{\ell}]$ has to be large for some $H$. However, for any input $H$ of interest, we can always run the algorithm and the returned information will be correct and certifiably so. As such, no a priori analysis of the convergence speed is needed in order to obtain rigorous approximations of the quantity of interest.
@@ Results
[…]
@@
We note that for finite systems, convergence happens after a finite number of steps, namely for $\ell$ such that $\Lambda_{\ell}=\Gamma$. While the convergence result becomes trivial in the finite case, the formulation of a convex program containing constraints for the Gibbs state still adds a novel tool for the numerical treatment of finite-sized quantum many-body systems. The more interesting case for our results is however that of infinite systems. Given a fixed computational budget, one can choose $\ell$ and run the corresponding program (3) and obtain a superset $\mathsf{p}^{\min}_{\ell},\mathsf{p}^{\max}_{\ell}]$ of the target interval $\langle O\rangle^{\min}_{\beta},\langle O\rangle^{\max}_{\beta}]$. In the cases where $\langle O\rangle^{\min}_{\beta}=\langle O\rangle^{\max}_{\beta}$, e.g., for finite systems at $T>0$, one can also run the program (3) with increasing values of $\ell$ until $\mathsf{p}^{\max}_{\ell}\leq\mathsf{p}^{\min}_{\ell}+\varepsilon$, where $\varepsilon$ is some desired accuracy $\varepsilon$. We note that, as previously mentioned, the problem of computing expectation values for equilibrium states is at least QMA-hard so we cannot hope to have fast convergence for all choices of $H$. But we stress that no a priori analysis of convergence speed is required to obtain some desired accuracy: as soon as we get $\mathsf{p}^{\max}_{\ell}\leq\mathsf{p}^{\min}_{\ell}+\varepsilon$, an additive $\varepsilon$ approximation is guaranteed for this instance. Numerical results illustrating this algorithm can be found in Section 4.
@@
[…]
@@
Note that this is a promise problem where the objective is to decide if all the equilibrium states have an expectation value per site >a>a or all the equilibrium states have an expectation value per site <a<a.
@@
Consider a translation-invariant local Hamiltonian $H$ on $\Gamma=\mathbb{Z}^{D}$ with $D\leq 2$. Assume furthermore that $H$ is commuting, i.e., $h_{X},h_{Y}]=0$ for all $X,Y\subset\mathbb{Z}^{D}$. For $D=1$ and $\beta_{1}=\infty$ or for $D=2$ and some $\beta_{1}>0$, we have for all $0\leq\beta<\beta_{1}$, and for any local observable $O$, $\langle O\rangle^{\min,\mathrm{TI}}_{\beta}=\langle O\rangle^{\max,\mathrm{TI}}_{\beta}=\langle O\rangle^{\mathrm{TI}}_{\beta}$ and, using the same notation as in Theorem 1.1
@@ 2 Preliminaries
where $h_{X}\in\mathcal{A}_{X}$ are the local interaction terms. We assume that the Hamiltonian $H$ is local, i.e., there exists a constant $r$ such that $h_{X}$ is nonzero only for $X$ having size at most $r$, for every site $x\in\Gamma$, $|\{X\subset\Gamma:x\in X,h_{X}\neq 0\}|\leq r$, and that $\|h_{X}\|\leq 1$ for all $X$. In the case of the infinite lattice $\Gamma=\mathbb{Z}^{D}$, we say that the Hamiltonian is translation-invariant if $h_{X+x}=\tau_{x}(h_{X})$ for all $X\subset\mathbb{Z}^{D}$ and $x\in\mathbb{Z}^{D}$, where $\tau_{x}:\mathcal{A}_{X}\to\mathcal{A}_{X+x}$ is the translation operator by $x$. If $\Lambda$ is a finite subset of $\Gamma$, we let
@@
where the blocks correspond to the ground-space and its orthogonal complement, we can see that $C$ has nonpositive diagonal entries, but is also positive semidefinite and thereby 0. By the Schur complement lemma we can then conclude that $B=0$ and thereby $\rho$ is supported on the ground space.
@@
[…]
For the finite Hamiltonian $\tilde{H}_{\Lambda}$ with $\Lambda=\{-\ell,\dots,\ell\}$, the unique ground state is the sequence $(2,0,\ldots,0)$ with energy 1-1. This makes it also the ground state of the infinite system corresponding to the limit with open boundary condition. Importantly, note that for the observable $O=|0\rangle\!\langle 0|$ at site 0, we have that for any $\ell\geq 1$, its value in the ground state of $H_{\Lambda}$ is $1$. However, another valid ground state of the infinite chain is the all-1 state. For this ground state in the thermodynamic limit, the observable takes a value of 0. It is realized by adding a boundary term $h_{\ell+1}(s)=-2\delta_{1,s}$ as it lowers the energy of the all-1 configuration below the open boundary ground state.
@@ 3.1 Energy-entropy balance inequalities
Let $H$ be a local Hamiltonian. Then $\omega:\mathcal{A}\to\mathbb{C}$ is a $\beta$-KMS state if, and only if, it satisfies [^2] The function $x\log(x/y)$ is defined as equal to 0 if $x=0$ and $\infty$ if $x>0$ and $y=0$.
@@
Similarly to the example for ground-states in Remark 2.1, we demonstrate how different thermal states can arise from different choices of boundary conditions. Consider the 2D Ising model at some temperature $0<T<T_{c}$ (where $T_{c}$ is the critical temperature), and let for any $\ell\in\mathbb{N}$, $\rho_{\ell,\beta}$ be the finite Gibbs state on a region $\{-\ell,\dots,\ell\}^{2}$ with open boundary condition. Then by the spin-flip symmetry of the Ising Hamiltonian, we know that $\operatorname{tr}(Z_{x}\rho_{\ell,\beta})=0$ for any site $x$ in the region $\{-\ell,\dots,\ell\}^{2}$. Thus this means that for any limit $\omega_{\beta}$ of the finite Gibbs state with open boundary condition, we will have $\omega_{\beta}(Z_{x})=0$. It is known however that for the 2D Ising model below the critical temperature, there are two (extremal) infinite-volume Gibbs states $\omega_{\beta}^{+}$ and $\omega_{\beta}^{-}$ such that $\omega_{\beta}^{-}(Z_{x})<0<\omega_{\beta}^{+}(Z_{x})$, which can be reached by taking limits of finite Hamiltonians with fixed boundary condition, i.e., aligned spins on the boundary [48].
@@ 3.2 Convex relaxations
The EEB inequalities above can be used to obtain rigorous lower and upper bounds on the value of any observable on thermal states via convex optimization. Let $(\Lambda_{\ell})_{\ell\in\mathbb{N}}$ be an increasing sequence of subsets of $\Gamma$ such that the support of $O$ is contained in $\Lambda_{0}$, and such that $\Lambda_{\ell}\uparrow\Gamma$ as $\ell\to\infty$. To simplify the presentation, we will further assume that $\overline{\Lambda_{\ell-1}}\subset\Lambda_{\ell}$ for all $\ell$, where, as before, $\overline{\Lambda}=\Lambda\cup\partial^{ex}\Lambda$ (see (21)). For example, if $\Gamma=\mathbb{Z}^{D}$ and $H$ is a nearest-neighbor Hamiltonian and $O$ is supported on site 0, one can take $\Lambda_{\ell}=\{-\ell,\ldots,\ell\}^{D}$. For any $\ell$, let $\mathcal{A}_{\ell}\subset\mathcal{A}$ be the subalgebra of observables supported on $\Lambda_{\ell}$. Consider the optimization problem
@@ 3.3 Convergence
$\langle O\rangle^{\min}_{\beta},\langle O\rangle^{\max}_{\beta}]=\{\omega(O):\omega\in\mathcal{G}_{\beta}\}.$ (48)
@@
so that we can write $\tilde{\omega}(b_{s}^{*}b_{t})]_{s,t}=K^{*}[\tilde{\omega}(a_{i}^{*}a_{j})]_{i,j}K$ and similarly for $\tilde{\omega}(b_{t}b_{s}^{*})]_{s,t}$. Hence we get
@@
For each $\ell$, let $\tilde{\omega}_{\ell}:\mathcal{A}_{\ell}\to\mathbb{C}$ be a feasible solution to (45) such that $\mathsf{p}^{\min}_{\ell}\leq\tilde{\omega}_{\ell}(O)\leq\mathsf{p}^{\min}_{\ell}+1/\ell$ (if the minimum value of the optimization problem is attained, we simply take $\tilde{\omega}_{\ell}$ the optimal solution which satisfies $\tilde{\omega}_{\ell}(O)=\mathsf{p}^{\min}_{\ell}$). We can extend $\tilde{\omega}_{\ell}$ to a state on the full $C^{*}$-algebra $\mathcal{A}$ [47, Proposition 2.3.24]. We now have a sequence of states $\tilde{\omega}_{\ell}$ on $\mathcal{A}$. Since the set of states on $\mathcal{A}$ is weak-* compact [47, Theorem 2.3.15], we can extract from $\tilde{\omega}_{\ell}$ a convergent subnet $\tilde{\omega}^{\prime}_{i}$ defined on the directed set $\mathcal{I}$ via a monotone final function $h:\mathcal{I}\to\mathbb{N}$, $\tilde{\omega}^{\prime}_{i}:=\tilde{\omega}_{h(i)}$. We denote the limit of the subnet by $\tilde{\omega}:=\lim_{i}\tilde{\omega}^{\prime}_{i}$, which is a state on $\mathcal{A}$.
We now show that $\tilde{\omega}$ is a $\beta$-KMS state by showing that the conditions (2) hold. Let $a\in\mathcal{A}_{\text{loc}}$. Since $a$ is a local observable we have for some $\ell^{\prime}$ that $a\in\mathcal{A}_{\ell}$ for all $\ell\geq\ell^{\prime}$. Hence this implies that
$\tilde{\omega}_{\ell}([H,a])=0$ (51)
$\tilde{\omega}_{\ell}(a^{*}a)\log\left(\frac{\tilde{\omega}_{\ell}(a^{*}a)}{\tilde{\omega}_{\ell}(aa^{*})}\right)\leq\beta\tilde{\omega}_{\ell}(a^{*}[H,a])$
for all $\ell\geq\ell^{\prime}$. Since $h$ is final and monotone, there exists $i^{\prime}\in\mathcal{I}$, such that $h(i)\geq\ell^{\prime}$ $\forall i\geq i^{\prime}$. Let $S$ be the complement of the set of states on which (LABEL:eq:EEBlimitCond) holds, which is open by weak-* continuity of $\omega\mapsto\omega(O)$ for any $O\in\mathcal{A}$. Assume $\tilde{\omega}\in S$. Then there exists $k\in\mathcal{I}$ such that $\forall i\geq k$ we have $\tilde{\omega}_{i}^{\prime}\in S$. However, since $\mathcal{I}$ is a directed set, there exists $i\in\mathcal{I}$ with $i\geq i^{\prime}$ and $i\geq k$ and so $\tilde{\omega}_{h(i)}\not\in S$ and $\tilde{\omega}_{h(i)}\in S$, which is a contradiction. Therefore we conclude that $\tilde{\omega}$ fulfills (LABEL:eq:EEBlimitCond) or likewise (2) and is a $\beta$-KMS state.
Finally note that
$\tilde{\omega}(O)=\lim_{i}\tilde{\omega}^{\prime}_{i}(O)=\lim_{i}\tilde{\omega}_{h(i)}(O)\leq\lim_{i}\mathsf{p}^{\min}_{h(i)}+1/h(i)=\mathsf{p}^{\min}_{\infty}.$ (52)
Here, $\mathsf{p}^{\min}_{h(i)}$ and $1/h(i)$ are subnets of the convergent sequences $\mathsf{p}^{\min}_{\ell}$ and $1/\ell$ and thereby converge. Since $\mathsf{p}^{\min}_{\infty}\leq\langle O\rangle^{\min}_{\beta}\leq\tilde{\omega}(O)$ we necessarily get equality $\mathsf{p}^{\min}_{\infty}=\langle O\rangle^{\min}_{\beta}$. The same argument can be used to show that $\mathsf{p}^{\max}_{\infty}=\langle O\rangle^{\max}_{\beta}$. ∎
@@ 3.4 Decidability
[…]
and similarly for $\tilde{\omega}(a_{j}a_{i}^{*})]_{ij}$. We use this to strengthen our convex relaxation (45) by adding the contraints
@@
After this modification, since the moment matrix is bounded away from zero and upper bounded by 1, the eigenvalues in the argument of the logarithm for the matrix relative entropy are contained in an interval $[a,1/a]$ for some $a>0$. We now introduce a semidefinite approximation of $D_{op}$ from [17]. For any integer $m\geq 1$ let $h_{m}(x)=2^{m}\left(x^{1/2^{m}}-1\right)$ which satisfy $\log(x)\leq h_{m+1}(x)\leq h_{m}(x)$ and note that $h_{m}\to\log$ uniformly on any compact interval in $(0,\infty)$ since $0\leq h_{m}(x)-\log(x)\leq 2^{-m}[(x-1)^{2}+(x^{-1}-1)^{2}]$ for all $x>0$. Consider the noncommutative perspective [51, 52] of $h_{m}$ which we denote by $D_{op}^{[m]}$:
@@
Then $D_{op}^{[m]}$ is jointly convex in $(X,Y)$, has a finite semidefinite programming formulation [17], and satisfies $D_{op}^{[m]}(X\|Y)\preceq D_{op}(X\|Y)$ since $h_{m}\leq-\log$. By changing the function $D_{op}$ in the entropy constraint (45e) by $D_{op}^{[m]}$ we get a semidefinite program $(\mathsf{P}^{{}^{\prime}\min}_{\ell,m})$ with rational coefficients whose optimal value satisfies $\mathsf{p}^{{}^{\prime}\min}_{\ell,m}\leq\mathsf{p}^{{}^{\prime}\min}_{\ell}$ and analogously for the upper bounds.